A force of 2N stretches an elastic material by 15mm. What additional force will stretch the material 17mm , assume that the elastic limit is not exceeded.
1
Expert's answer
2020-07-02T17:14:43-0400

Find the stiffness of the material from the given data: 2 N stretch it for 15 mm:


"k=\frac{F}{x}=\frac{2}{0.015}=133.3\text{ N/m}."

For 17 mm, the force required is


"F_1=kx_1=133.3\cdot0.017=2.267\text{ N}."

The additional force, therefore, is


"\Delta F=F_1-F=0.267\text{ N}."